OpenStudy (anonymous):

a rock falls from a tower that is 432 feet high. as it is falling, its height is given by the formula h=432-16ft^2. how man secods will it take for the rock to hit the ground (h=0)

OpenStudy (anonymous):

you mean? \[h=432-16t^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

put 0 instead h (as it said) and solve quadratic equation

OpenStudy (anonymous):

20.4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how you got 20.4?

OpenStudy (anonymous):

mental math

OpenStudy (anonymous):

then do it on paper and you will get correct answer :D

OpenStudy (anonymous):

didn't serve me so well

OpenStudy (lalaly):

0 = 432 - 16t² 0 = 16(27 - t²) 27 - t² = 0 -t² = -27 t² = 27 t = 3√3

OpenStudy (anonymous):

oh you don't need to do that because you got answer :/

OpenStudy (anonymous):

not technically. i needed 5.2 seconds

OpenStudy (anonymous):

y^2=18 9 correct?

OpenStudy (anonymous):

i think you should write it precisely 3sqrt3 unless told otherwise

OpenStudy (anonymous):

no, 9*9=81

OpenStudy (anonymous):

was told otherwise, to write in seconds.

OpenStudy (anonymous):

you can easily write \(3\sqrt3\) second

OpenStudy (anonymous):

ya, 5.2

OpenStudy (anonymous):

\[\sqrt{18}=\sqrt{2\cdot9}=3\sqrt2\]

OpenStudy (anonymous):

well write how you want, idk how you are being taught in your country

OpenStudy (anonymous):

another problem i can't figure is y^2=18

OpenStudy (anonymous):

not very well, i am 24 and don't get this junk

OpenStudy (anonymous):

\[\sqrt{18}=\sqrt{2\cdot9}=3\sqrt2\]

OpenStudy (anonymous):

not an answer

OpenStudy (anonymous):

then your answers are incorrect

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sqrt18

OpenStudy (anonymous):

324, +,- sqrt18, 9, or sqrt18

OpenStudy (anonymous):

i did, doesn't work either